OpenStudy (anonymous):

If y = 5x – 4, which of the following sets represents possible inputs and outputs of the function, represented as ordered pairs? (4 points) {(0, 4), (2, –6), (4, 20)} {(0, –4), (2, 6), (4, 16)} {(0, –4), (2, 6), (4, 20)} {(0, 4), (2, 6), (4, 16)}

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

Each coordinate is \((x,y)\). You want plug in the \(x\) values \(\{0,2,4\}\) to figure out the corresponding \(y\) values.

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

For example, for \(x=0\):\[ (x, y)=(x,5x-4) = (0,5(0)-4) = (0,-4) \]So one point is \((0,-4)\).

OpenStudy (anonymous):

oh! wait..... i dont get it....

OpenStudy (anonymous):

This eliminates all but: \[ \{(0, -4), (2, 6), (4, 16)\}\\ \{(0,-4), (2, 6), (4, 20)\} \]

OpenStudy (anonymous):

Okay, first step: We know that \(y=5x-4\), right? This means wherever there is a \(y\), we can put \(5x-4\) and it will still be equal. So: \[ (x,y) = (x,5x-4) \]Does this make sense?

OpenStudy (anonymous):

okay so it would be the 2nd one?

OpenStudy (anonymous):

right?!

OpenStudy (anonymous):

Do you mean \[ \{(0,-4),(2,6),(4,16)\} \]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is that right?!

OpenStudy (anonymous):

Yes, it is correct.